OpenStudy (anonymous):

The equation of a circle is shown below. Determine the radius. (x − 1)2 + (y + 1)2 = 121 Select one: A. 9 B. 10 C. 11 D. 12

OpenStudy (anonymous):

compare it with (x-x1)^2+(y-y1^2=c^2 herre c=radius...

OpenStudy (anonymous):

Its 11!!!!

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] (h,k) are the center of the circle and r is the radius but as you see it is squared so you can take the square root of that value to find the actual radius.